OpenStudy (anonymous):

how to find perimeter of: 4√3 + 9

OpenStudy (anonymous):

is there anyone to help out or not?

OpenStudy (jdoe0001):

based on the definition of a perimeter... that value above has none to begin with a perimeter has no relation unless you give it specific context.....

OpenStudy (anonymous):

i don't get it.. can you solve it?

OpenStudy (jdoe0001):

do you know what a perimeter is?

OpenStudy (anonymous):

yeah i do but i just don't know how to do it with square roots..

OpenStudy (jdoe0001):

hmm so you just want the SUM of those 2 numbers then ... well, if you want a decimal figure, just use the calculator if not, then leave it like it's :)

OpenStudy (anonymous):

what?? never mind.. thanks!!

OpenStudy (jdoe0001):

\(\bf \sqrt{3} \approx 1.73\qquad 4\sqrt{3}=4\cdot 1.73\qquad 4\sqrt{3}+9=4\cdot 1.73+9\)

OpenStudy (anonymous):

sorry i don't get it..

OpenStudy (anonymous):

can you just tell the answer?

OpenStudy (jdoe0001):

well, you just have the sum of 2 numbers numbers do not have a perimeter, shapes or figures do

OpenStudy (anonymous):

how come?

OpenStudy (jdoe0001):

well.. in what context are you defining "perimeter" in this instance?

OpenStudy (anonymous):

if i know that then i can solve it by myself..

OpenStudy (anonymous):

this question just says to find area and perimeter and i do find area but just confuse in finding perimeter..

OpenStudy (jdoe0001):

ok... that means you have a figure.... well, can you take a quick screenshot of the material, so we can see the figure?

OpenStudy (anonymous):

no it doesn't have any figure.. its an statement question..

OpenStudy (jdoe0001):

well, is very ambiguous then, maybe someone else may know

OpenStudy (anonymous):

i don't think so but thanks for trying..